# Hyperbolic Busemann Neural Networks

### v1

@@ 4.2 Generalization
This is reflected in Thms.˜4.2 and 4.3 by replacing every $u_{k}(x)$ with $\phi\left(-\alpha_{k}B^{v_{k}}(x)+b_{k}\right)$. Moreover, inspired by the Poincaré Möbius transformation [21, Sec. 3.2], a BFC transformation could be further followed by a gyroaddition $\oplus_{\mathcal{H}}$: $\mathcal{H}^{n}_{K}\ni x\mapsto\mathcal{F}(x)\oplus_{\mathcal{H}}b\in\mathcal{H}^{m}_{K}$ with $b\in\mathcal{H}^{m}_{K}$ as a gyro bias. For example, the Lorentz BFC layer is generalized as
@@ 5.1 Image classification
Setup. Following Bdeir et al. [4, Sec. 5.1] and Guo et al. [26, Sec. 4], we use a hybrid architecture with a ResNet-18 [27] backbone and an MLR head. We compare Euclidean MLR with hyperbolic variants in both models. In Poincaré, we evaluate Poincaré MLR (PMLR) re-parameterized by Shimizu et al. [49], Pseudo Busemann MLR (PBMLR-P) [42], and our BMLR-P. In Lorentz, we evaluate Lorentz MLR (LMLR) [4] and our BMLR-L. For hyperbolic MLRs, we map the ResNet-18 features to the target hyperbolic space before classification. We evaluate on CIFAR-10 [34], CIFAR-100 [34], Tiny-ImageNet [35], and ImageNet-1k [15]. On the first three datasets, we conduct five-fold experiments. More details appear in Sec.˜E.1.
@@ B.3.1 General definition
Classical vector spaces can be characterized as a commutative group together with a compatible scalar multiplication. By analogy, a gyrovector space is built from a gyrocommutative gyrogroup endowed with a compatible scalar gyromultiplication [53, 43, 13].
@@
Definition B.15 (Nonreductive gyrogroup [43]).
A groupoid $(G,\oplus)$ is a nonreductive gyrogroup if it satisfies axioms (G1), (G2), and (G3).
@@
Remark B.17.
Nguyen [43, Def. 2.3] presented a related definition in which (V1) additionally states $0\odot x=t\odot e=e$ and $(-1)\odot x=\ominus x$. By Ungar [53, Thm. 6.4], these statements follow from the axioms above and are therefore redundant. Ungar [53, Def. 6.2] also introduced real inner product gyrovector spaces by imposing extra inner product axioms, which are not required here.
@@
B.3.2 Gyro-structures over manifolds
Following Nguyen [43], for $x,y,z\in\mathcal{M}$ and $t\in\mathbb{R}$, the gyro-structure on the manifold $\mathcal{M}$ is defined by
[…]
Here $e\in\mathcal{M}$ is a pre-defined identity element. These definitions are valid once the axioms of a gyrovector space hold. As demonstrated by Chen et al. [13, Sec. 5], this construction equips several manifolds, including the Poincaré and Lorentz models, with gyrovector spaces.
@@ B.4 Hyperbolic geometry
where $e$ denotes the origin in $\mathcal{H}^{n}_{K}$. On the Poincaré ball $\mathbb{P}^{n}_{K}$, such gyro-structure is known as the Möbius gyrovector space [53, Ch. 6.14]:
@@
where $\ominus_{\mathrm{M}}x=-1\odot_{\mathrm{M}}x=-x$ is the gyroinverse and $\mathbf{0}$ is the gyro identity: $\mathbf{0}\oplus_{\mathrm{M}}x=x,\forall x\in\mathbb{P}^{n}_{K}$. As shown by Chen et al. [13, Props. 24-25], the Lorentz gyroaddition and gyromultiplication also admit closed-form expressions:
@@ Appendix C Comparison with existing hyperbolic MLR
Poincaré MLR [21] Geodesic [21, Def. 3.1] $\left\{x\in\mathbb{P}^{n}_{K}\mid\left\langle\operatorname{Log}_{p}\left(x\right),a\right\rangle_{p}=0\right\}$ $p\in\mathbb{P}^{n}_{K},\ a\in T_{p}\mathbb{P}^{n}_{K}$ $\mathbb{P}^{n}_{K}$
Pseudo-Busemann MLR [42] Busemann & gyro [42, Def. 4.1] $\left\{x\in\mathbb{P}^{n}_{K}\mid B^{v}\left(-p\oplus_{\mathrm{M}}x\right)=0\right\}$ $v\in\mathbb{S}^{n-1},\ p\in\mathbb{P}^{n}_{K}$ $\mathbb{P}^{n}_{K}$
@@
Poincaré MLR [21] $\frac{1}{\sqrt{-K}}\sinh^{-1}\left(\frac{2\sqrt{-K}\left|\left\langle-p\oplus_{\mathrm{M}}x,a\right\rangle\right|}{\left(1+K\left\|-p\oplus_{\mathrm{M}}x\right\|^{2}\right)\left\|a\right\|}\right)$ [21, Thm. 5] $\mathbb{P}^{n}_{K}$ Real
Pseudo-Busemann MLR [42] $\operatorname{d}(x,p)\frac{B^{v}\left(-p\oplus_{\mathrm{M}}x\right)}{\left\|-p\oplus_{\mathrm{M}}x\right\|}$ [42, Cor. 4.3] $\mathbb{P}^{n}_{K}$ Pseudo
@@
In addition to Tab.˜2, Tabs.˜11 and 12 further compare hyperplanes and point-to-hyperplane distances across methods. We draw the following three conclusions.
@@ D.1 Poincaré model
Using Eq.˜3 with $v=e_{k}$, Eq.˜47 becomes
@@
Define $t_{k}=\exp\left(-\sqrt{-K}u_{k}(x)\right)>0$. Exponentiating Eq.˜48 gives
@@
Writing $R=\left\|y\right\|^{2}$, Eq.˜49 yields an affine expression for each coordinate
@@
A real solution $R$ exists only if $\Delta_{\mathrm{P}}\geq 0$. In addition, feasibility requires $0\leq R<-1/K$ so that $y\in\mathbb{P}^{m}_{K}$. Such conditions can fail for generic $\{u_{k}(x)\}$, in which case no $y\in\mathbb{P}^{m}_{K}$ satisfies Eq.˜47.
@@ D.2 Lorentz model
Using Eq.˜4 with $v=e_{k}$, Eq.˜47 becomes
@@
Define $t_{k}=\exp\left(-\sqrt{-K}u_{k}(x)\right)>0$ and, denoting $T=\sum_{k=1}^{m}t_{k}$ and $q=\sum_{k=1}^{m}t_{k}^{2}$, we obtain from Eq.˜54,
@@
Hence, a real $y_{t}$ exists only if $T^{2}-(m-1)\left(1+q\right)\geq 0$. In addition, feasibility requires $y_{t}>0$. Such conditions can fail for generic $\{u_{k}(x)\}$, where no $y\in\mathbb{L}^{m}_{K}$ satisfies Eq.˜47.
@@ D.3 Summary
Equating Busemann coordinates as in Eq.˜47 requires nontrivial inequalities on the responses $\{u_{k}(x)\}$. These constraints are not guaranteed during learning, so the system can become infeasible and the output $y$ undefined. This motivates our choice in Eq.˜22 to use signed point-to-hyperplane distances on the LHS, which admit closed-form solutions that are feasible for all inputs and parameters in both the Poincaré and Lorentz models.
@@ E.1.2 Implementation details
For CIFAR-10/100 and Tiny-ImageNet, we follow Bdeir et al. [4, App. C.1]. For ImageNet-1k, we follow Guo et al. [26, Sec. 4]. Tab.˜13 summarizes the dataset-specific hyperparameters. For the hyperbolic MLR, before mapping into the hyperbolic space, we clip the feature vector by
@@
All methods are implemented in PyTorch and trained with cross-entropy loss. The results of MLR, PMLR, and LMLR on CIFAR-10/100 and Tiny-ImageNet are copied from Bdeir et al. [4, Tab. 1], while the ones of PBMLR-P on CIFAR-10/100 are copied from Nguyen et al. [42, Tab. 2]. The remaining results are obtained by our careful implementation.
@@ E.2.2 Implementation details
We mainly follow the official implementations of Khan et al. [31] for data processing, model architecture, and training. We adopt a simple CNN with three convolutional blocks followed by dense, ReLU activated layers to extract features [31, Fig. 4]. Before classification, the features are clipped and mapped to the target hyperbolic space as in Eq.˜58, then passed to either a prior hyperbolic MLR or our BMLR head. The clipping factor defaults to $r=1$, with the following exceptions on GUE: on the Lorentz model, $r=2.0$ for Covid variant classification and $r=5.0$ for species classification; on the Poincaré ball, $r=2.0$ for species classification. Following Khan et al. [31], we treat the curvature $K$ as a learnable parameter initialized as 1-1. The remaining hyperparameters are listed in Tab.˜16. All models share these hyperparameters, except LMLR on Covid variant classification, for which we set the weight decay to $1e^{-3}$ to ensure convergence.
@@ E.3.2 Implementation details
All methods are implemented in PyTorch and run on a single RTX A6000 GPU. The results of HGCN-PMLR and HGCN-PBMLR-P on the Poincaré ball are taken from Nguyen et al. [42, Tab. 11]. Results for the remaining baselines are obtained from our reimplementation following the original settings.
@@ Appendix F Proofs
Taking infimum in Eq.˜67 over $y\in H^{\gamma}_{\tau_{1}}$ gives
@@
Step 4: sandwich closure. Combining Eqs.˜68 and 70 gives
@@
The hyperplane and point-to-hyperplane distance are presented in Tabs.˜11 and 12. The origin of the Poincaré ball model is $e=\mathbf{0}\in\mathbb{P}^{m}_{K}$. The specific ones w.r.t. the origin [49, Def. 1 and Eq. (56)] are
@@
Note that Eq.˜77 takes the same form as Shimizu et al. [49, Eq. (56)], except that their responses are different. The below proof are inspired by their derivation.
Applying $\sinh(\cdot)$ on both sides of Eq.˜77 yields
@@
Hence, $y$ is collinear with $\omega$. Write $y=\lambda\omega$ with $\lambda\geq 0$. Substituting into Eq.˜80 and taking norms gives a quadratic in $\lambda$:
@@
Applying $\sinh(\cdot)$ to both sides of Eq.˜88 yields

### v2

@@ 4.2 Generalization
This is reflected in Thms.˜4.2 and 4.3 by replacing every $u_{k}(x)$ with $\phi\left(-\alpha_{k}B^{v_{k}}(x)+b_{k}\right)$. Moreover, inspired by the Poincaré Möbius transformation [23, Sec. 3.2], a BFC transformation could be further followed by a gyroaddition $\oplus_{\mathcal{H}}$: $\mathcal{H}^{n}_{K}\ni x\mapsto\mathcal{F}(x)\oplus_{\mathcal{H}}b\in\mathcal{H}^{m}_{K}$ with $b\in\mathcal{H}^{m}_{K}$ as a gyro bias. For example, the Lorentz BFC layer is generalized as
@@ 5.1 Image classification
Setup. Following Bdeir et al. [4, Sec. 5.1] and Guo et al. [28, Sec. 4], we use a hybrid architecture with a ResNet-18 [29] backbone and an MLR head. We compare Euclidean MLR with hyperbolic variants in both models. In Poincaré, we evaluate Poincaré MLR (PMLR) re-parameterized by Shimizu et al. [51], Pseudo Busemann MLR (PBMLR-P) [45], and our BMLR-P. In Lorentz, we evaluate Lorentz MLR (LMLR) [4] and our BMLR-L. For hyperbolic MLRs, we map the ResNet-18 features to the target hyperbolic space before classification. We evaluate on CIFAR-10 [36], CIFAR-100 [36], Tiny-ImageNet [37], and ImageNet-1k [17]. On the first three datasets, we conduct five-fold experiments. More details appear in Sec.˜E.1.
@@ B.3.1 General definition
Classical vector spaces can be characterized as a commutative group together with a compatible scalar multiplication. By analogy, a gyrovector space is built from a gyrocommutative gyrogroup endowed with a compatible scalar gyromultiplication [55].
@@ B.4 Hyperbolic geometry
where $e$ denotes the origin in $\mathcal{H}^{n}_{K}$. On the Poincaré ball $\mathbb{P}^{n}_{K}$, such gyro-structure is known as the Möbius gyrovector space [55, Ch. 6.14]:
@@
where $\ominus_{\mathrm{M}}x=-1\odot_{\mathrm{M}}x=-x$ is the gyroinverse and $\mathbf{0}$ is the gyro identity: $\mathbf{0}\oplus_{\mathrm{M}}x=x,\forall x\in\mathbb{P}^{n}_{K}$. As shown by Chen et al. [15, Props. 24-25], the Lorentz gyroaddition and gyromultiplication also admit closed-form expressions:
@@ Appendix C Comparison with existing hyperbolic MLR
Poincaré MLR [23] Geodesic [23, Def. 3.1] $\left\{x\in\mathbb{P}^{n}_{K}\mid\left\langle\operatorname{Log}_{p}\left(x\right),a\right\rangle_{p}=0\right\}$ $p\in\mathbb{P}^{n}_{K},\ a\in T_{p}\mathbb{P}^{n}_{K}$ $\mathbb{P}^{n}_{K}$
Pseudo-Busemann MLR [45] Busemann & gyro [45, Def. 4.1] $\left\{x\in\mathbb{P}^{n}_{K}\mid B^{v}\left(-p\oplus_{\mathrm{M}}x\right)=0\right\}$ $v\in\mathbb{S}^{n-1},\ p\in\mathbb{P}^{n}_{K}$ $\mathbb{P}^{n}_{K}$
@@
Poincaré MLR [23] $\frac{1}{\sqrt{-K}}\sinh^{-1}\left(\frac{2\sqrt{-K}\left|\left\langle-p\oplus_{\mathrm{M}}x,a\right\rangle\right|}{\left(1+K\left\|-p\oplus_{\mathrm{M}}x\right\|^{2}\right)\left\|a\right\|}\right)$ [23, Thm. 5] $\mathbb{P}^{n}_{K}$ Real
Pseudo-Busemann MLR [45] $\operatorname{d}(x,p)\frac{B^{v}\left(-p\oplus_{\mathrm{M}}x\right)}{\left\|-p\oplus_{\mathrm{M}}x\right\|}$ [45, Cor. 4.3] $\mathbb{P}^{n}_{K}$ Pseudo
@@
Hyperbolic MLRs follow a point-to-hyperplane formulation, of which Chen et al. [12, Eqs. 4-6] provides the Riemannian prototype. Therefore, the key difference lies in hyperplanes and point-to-hyperplane distances across methods. In addition to Tab.˜2, Tabs.˜11 and 12 further make this comparison. We draw the following three conclusions.
@@ D.1 Poincaré model
Using Eq.˜3 with $v=e_{k}$, Eq.˜43 becomes
@@
Define $t_{k}=\exp\left(-\sqrt{-K}u_{k}(x)\right)>0$. Exponentiating Eq.˜44 gives
@@
Writing $R=\left\|y\right\|^{2}$, Eq.˜45 yields an affine expression for each coordinate
@@
A real solution $R$ exists only if $\Delta_{\mathrm{P}}\geq 0$. In addition, feasibility requires $0\leq R<-1/K$ so that $y\in\mathbb{P}^{m}_{K}$. Such conditions can fail for generic $\{u_{k}(x)\}$, in which case no $y\in\mathbb{P}^{m}_{K}$ satisfies Eq.˜43.
@@ D.2 Lorentz model
Using Eq.˜4 with $v=e_{k}$, Eq.˜43 becomes
@@
Define $t_{k}=\exp\left(-\sqrt{-K}u_{k}(x)\right)>0$ and, denoting $T=\sum_{k=1}^{m}t_{k}$ and $q=\sum_{k=1}^{m}t_{k}^{2}$, we obtain from Eq.˜50,
@@
Hence, a real $y_{t}$ exists only if $T^{2}-(m-1)\left(1+q\right)\geq 0$. In addition, feasibility requires $y_{t}>0$. Such conditions can fail for generic $\{u_{k}(x)\}$, where no $y\in\mathbb{L}^{m}_{K}$ satisfies Eq.˜43.
@@ D.3 Summary
Equating Busemann coordinates as in Eq.˜43 requires nontrivial inequalities on the responses $\{u_{k}(x)\}$. These constraints are not guaranteed during learning, so the system can become infeasible and the output $y$ undefined. This motivates our choice in Eq.˜22 to use signed point-to-hyperplane distances on the LHS, which admit closed-form solutions that are feasible for all inputs and parameters in both the Poincaré and Lorentz models.
@@ E.1.2 Implementation details
For CIFAR-10/100 and Tiny-ImageNet, we follow Bdeir et al. [4, App. C.1]. For ImageNet-1k, we follow Guo et al. [28, Sec. 4]. Tab.˜13 summarizes the dataset-specific hyperparameters. For the hyperbolic MLR, before mapping into the hyperbolic space, we clip the feature vector by
@@
All methods are implemented in PyTorch and trained with cross-entropy loss. The results of MLR, PMLR, and LMLR on CIFAR-10/100 and Tiny-ImageNet are copied from Bdeir et al. [4, Tab. 1], while the ones of PBMLR-P on CIFAR-10/100 are copied from Nguyen et al. [45, Tab. 2]. The remaining results are obtained by our careful implementation.
@@ E.2.2 Implementation details
We mainly follow the official implementations of Khan et al. [33] for data processing, model architecture, and training. We adopt a simple CNN with three convolutional blocks followed by dense, ReLU activated layers to extract features [33, Fig. 4]. Before classification, the features are clipped and mapped to the target hyperbolic space as in Eq.˜54, then passed to either a prior hyperbolic MLR or our BMLR head. The clipping factor defaults to $r=1$, with the following exceptions on GUE: on the Lorentz model, $r=2.0$ for Covid variant classification and $r=5.0$ for species classification; on the Poincaré ball, $r=2.0$ for species classification. Following Khan et al. [33], we treat the curvature $K$ as a learnable parameter initialized as 1-1. The remaining hyperparameters are listed in Tab.˜16. All models share these hyperparameters, except LMLR on Covid variant classification, for which we set the weight decay to $1e^{-3}$ to ensure convergence.
@@ E.3.2 Implementation details
All methods are implemented in PyTorch and run on a single RTX A6000 GPU. The results of HGCN-PMLR and HGCN-PBMLR-P on the Poincaré ball are taken from Nguyen et al. [45, Tab. 11]. Results for the remaining baselines are obtained from our reimplementation following the original settings.
@@ Appendix F Proofs
Taking infimum in Eq.˜63 over $y\in H^{\gamma}_{\tau_{1}}$ gives
@@
Step 4: sandwich closure. Combining Eqs.˜64 and 66 gives
@@
The hyperplane and point-to-hyperplane distance are presented in Tabs.˜11 and 12. The origin of the Poincaré ball model is $e=\mathbf{0}\in\mathbb{P}^{m}_{K}$. The specific ones w.r.t. the origin [51, Def. 1 and Eq. (56)] are
@@
Note that Eq.˜73 takes the same form as Shimizu et al. [51, Eq. (56)], except that their responses are different. The below proof are inspired by their derivation.
Applying $\sinh(\cdot)$ on both sides of Eq.˜73 yields
@@
Hence, $y$ is collinear with $\omega$. Write $y=\lambda\omega$ with $\lambda\geq 0$. Substituting into Eq.˜76 and taking norms gives a quadratic in $\lambda$:
@@
Applying $\sinh(\cdot)$ to both sides of Eq.˜84 yields